OpenStudy (anonymous):

Which values of x and y are solutions to the open sentence? 3y ≥ 2x + 5 A. x = 1, y = 1 B. x = 1, y = 2 C. x = 2, y = 2 D. x = 2, y = 3

OpenStudy (anonymous):



OpenStudy (welshfella):

that means 'greater or equal to'

OpenStudy (anonymous):

like 4 and 3

OpenStudy (anonymous):

i dont get the equal part

OpenStudy (welshfella):

x >= 2 that means x can be 2 or any number greater than 2

OpenStudy (anonymous):

oh thank you

OpenStudy (welshfella):

yw for the first part plug in values of x and y given in the options and see which fit 3y >= 2x + 5

OpenStudy (welshfella):

for example A x=1,y = 1 :- 3(1) >= 2(1) + 5 3 >= 7 which is not true so it cant be A

OpenStudy (anonymous):

for b 3 times 2= 6 and 2 times 1=2+5=7

OpenStudy (welshfella):

yes so b is not the correct option

OpenStudy (welshfella):

try d

OpenStudy (anonymous):

for d 3 times 3=9 and 2 times 2=4+5=9 i think d is right

OpenStudy (welshfella):

yes its D